# Higher-dimensional magnetic Skyrmions

### v1

@@ Abstract
We propose a generalization of the theory of magnetic Skyrmions in chiral magnets in two dimensions to a higher-dimensional theory with magnetic Skyrmions in three dimensions and an $S^{3}$ target space, requiring a 4-dimensional magnetization vector. A physical realization of our theory necessitates the use of a synthetic dimension, recently promoted and realized in condensed matter physics. In the simplest incarnation of the theory, we find a Skyrmion and a sphaleron – the latter being an unstable soliton. Including also the Skyrme term in theory enriches the spectrum to a small metastable Skyrmion, an unstable sphaleron and a large stable Skyrmion.
@@ 1 Introduction
The magnetic Skyrmion is gaining recognition as a prominent soliton in 2-dimensional magnetic physics, particularly in the context of chiral magnets and thin films [1, 2, 3, 4, 5], as well as spintronics and quantum technologies [6, 7, 8, 9]. The name Skyrmion comes from the 3-dimensional soliton in the Skyrme model, proposed by T. Skyrme in the 1960s as a model of the nucleus [10, 11], further validated as a low-energy effective field theory of QCD at large $N_{c}$ (number of colors) by E. Witten [12, 13] and later used as a validation of the Sakai-Sugimoto model as evidence that the model really describes QCD at low energies [14]. The Skyrmion in 2 or 3 dimensions is a texture soliton, which differs from a defect-type soliton by mapping the entire 2- or 3-dimensional space to the target space. In particular, a magnetic Skyrmion in two dimensions is mapping the entire plane to an $S^{2}$ target space, with the latter being the normalized magnetization vector $\bm{n}=(n^{1},n^{2},n^{3})$ such that $\bm{n}\cdot\bm{n}=1$ [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. It differs drastically from the 3-dimensional Skyrmion by the fact that: 1) it is stabilized by a lower-dimensional operator (i.e. less derivatives than number of spatial dimensions) and 2) it is chiral. The latter implies that magnetic Skyrmions stably exist, but no anti-Skyrmions exist. This is different from 3-dimensional Skyrmions, where a higher-dimensional operator (i.e. more derivatives than number of spatial dimensions) is used to stabilize the soliton – called a Skyrmion – against collapse. Indeed the higher-derivative (or smaller-derivative) terms are a way to avoid Derrick’s theorem [29].
@@
In this paper, we consider the generalization of the 2-dimensional magnetic Skyrmion with target space $S^{2}$ to a 3-dimensional magnetic Skyrmion with target space $S^{3}$. [^1] A preliminary investigation started in [30]. Since the unit magnetization vector in three dimensions gives rise to $S^{2}$, we need a unit magnetization vector in four dimensions. [^2] We use the notation “magnetization vector” for the 4-dimensional generalization of the 3-dimension magnetization vector, although it is physically different. For example, it is not possible to write $\bm{B}\cdot\bm{n}$, since the 3-dimensional magnetic field has 3 components and a would-be 4-dimensional magnetic field cannot be dualized to a vector (the 4-dimensional Hodge dual of a 2-form is also a 2-form). We will henceforth denote the construct as a higher-dimensional magnetic Skyrmion.
@@
The physical realization of a 3-dimensional magnetic Skyrmion would necessitate the incorporation of a synthetic (extra) dimension, as has been demonstrated in a number of physical systems – e.g. see Refs. [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Until now, the 3-dimensional counterpart of 2D magnetic Skyrmions were magnetic Hopfions, 3-dimensional solitons living in a magnetization vector that is also living in three dimensions. Recent numerical computations [41, 42, 43, 44, 45] suggest that similar objects may exist at the nanoscale in frustrated and chiral magnets. Micrometer-sized Hopfions have recently been created and observed experimentally in chiral ferromagnetic fluids [46, 47]. The 3D magnetic Skyrmions which we propose, instead, have $S^{3}$ as a target space, the same as used in the standard Skyrme model in three dimensions.
@@
In this paper, we concentrate on the spherically symmetric case which enjoys $\operatorname{SO}(3)$ invariance. We take the generalized Dzyaloshinskii-Moriya interaction (DMI) [48, 49, 50] to have only one derivative of the magnetization vector and a single component of the same vector. In four dimensions, this means that such a term must be contracted with a tensor that possesses two $\operatorname{SO}(4)$-indices and a single spatial index: we denote such a tensor by $\Theta$. $\operatorname{SO}(4)$ invariance dictates that $\Theta$ is antisymmetric in $\operatorname{SO}(4)$-indices and we use this fact, as well as global rotations to rotate the tensor to its standard form. We further impose $\operatorname{SO}(3)$ invariance, where we are locking spatial rotations with an $\operatorname{SO}(3)\subset\operatorname{SO}(4)$ rotation, in order to obtain a spherically symmetric soliton. This procedure leads to two different invariant structures in the $\Theta$ tensor. However, only one of these two structures remains nonvanishing when a spherically symmetric Ansatz is imposed. We furthermore show that both the Bloch-type and the Néel-type DMI generalizations lead to the same term when $\operatorname{SO}(3)$ invariance is imposed.
@@ 2 Higher-dimensional magnetic Skyrmions
A possible extension of planar magnetic Skyrmions can be realized by increasing the dimension of the base space and target space from two to three dimensions [30]. This amounts to a magnetization unit vector that would live in a 4-dimensional Cartesian space. For a physical realization, advanced concepts as synthetic dimensions must be considered, we will deal with this argument in Sec. 4. In symbols, the magnetization field is now a map $\bm{n}:\mathbb{R}^{3}\to S^{3}$ where $\bm{n}=(n^{1},n^{2},n^{3},n^{4})$ and $\bm{n}\cdot\bm{n}=1$. Let us consider the energy functional of the following form:
@@ 2.1.1 Bloch-type DMI
In order to put the higher-dimensional DM term on the most general standard form, we use $\operatorname{SO}(4)$ and $\operatorname{SO}(3)$ transformations to simplify the tensor $\Theta$ as much as possible. Since $\Theta$ is contracted with the $\operatorname{SO}(4)$-invariant tensor $\epsilon_{abcd}$, we can from here on assume that it is antisymmetric in the $\operatorname{SO}(4)$-indices $a$ and $b$:
@@
Performing simultaneously both orthogonal transformations, $R\in\operatorname{SO}(4)$ and $r\in\operatorname{SO}(3)$, we get
@@
which is invariant only under the $\operatorname{SO}(3)$ subgroup represented by $\rho\in\operatorname{SO}(3)\subset\operatorname{SO}(4)$ such that
@@
the rotation $\rho$ of Eq. (2.14) can only be used to diagonalize the submatrix consisting of $A^{i}$, $B^{i}$ and $C^{i}$. Since the submatrix is $3\times 3$, it only possesses 1 purely imaginary pair of nonvanishing eigenvalues, as well as a zero eigenvalue. There is still no reason for the $\operatorname{SO}(3)$ vectors to be aligned, so we can only diagonalize one of the $\Theta^{i}$ submatrices, say the $i=3$ again. We thus arrive at
$\Theta^{\alpha}=\frac{1}{2}\begin{pmatrix}0&A^{\alpha}&B^{\alpha}&D^{\alpha}\\ -A^{\alpha}&0&C^{\alpha}&E^{\alpha}\\ -B^{\alpha}&-C^{\alpha}&0&F^{\alpha}\\ -D^{\alpha}&-E^{\alpha}&-F^{\alpha}&0\end{pmatrix},\qquad\Theta^{3}=\frac{1}{2}\begin{pmatrix}0&A^{3}&0&D^{3}\\ -A^{3}&0&0&E^{3}\\ 0&0&0&F^{3}\\ -D^{3}&-E^{3}&-F^{3}&0\end{pmatrix},$ (2.16)
with $\alpha=1,2$. Using the $r\in\operatorname{SO}(3)$ degrees of freedom, we can eliminate one symbol in $\Theta^{\alpha}$, by rotating it into the $i=3$ direction; we may choose this to be $F^{\alpha}=0$. We will denote this the standard form of the $\Theta$ tensor. One result of this computation is that the minimal degrees of freedom of the $\Theta$ tensor are 14.
At this point, it will be useful to identify a subset of the $\Theta$ tensor in standard form, that is invariant under a combined $\operatorname{SO}(3)\subset\operatorname{SO}(4)$ and $\operatorname{SO}(3)$ symmetry. The transformation is again that of Eq. (2.10) with $R$ of Eq. (2.14), it takes the form
$\Theta^{i}\to\frac{1}{2}\begin{pmatrix}\rho^{\rm T}\begin{pmatrix}0&A^{j}&B^{j}\\ -A^{j}&0&C^{j}\\ -B^{j}&-C^{j}&0\end{pmatrix}\rho&\rho^{\rm T}\begin{pmatrix}D^{j}\\ E^{j}\\ F^{j}\end{pmatrix}\\ \begin{pmatrix}-D^{j}&-E^{j}&-F^{j}\end{pmatrix}\rho&0\end{pmatrix}r_{ji},$ (2.17)
where both $\rho$ and $r$ belongs to $\operatorname{SO}(3)$. From the above equation, it is clear that invariance requires the triplet $A,B,C$ to transform under the adjoint representation of $\operatorname{SO}(3)$ and the triplet $D,E,F$ to transform under the fundamental representation. This means that invariance under the combined rotation with the locking of the two $\operatorname{SO}(3)$’s according to Eq. (2.14) with $\rho=r$, is attained by the tensor
@@
$B^{\alpha}$ $=(-\Gamma^{3},0),$
$C^{\alpha}$ $=(0,-\Gamma^{3}),$
$D^{\alpha}$ $=(-\Gamma^{4},0),$
$E^{\alpha}$ $=(0,-\Gamma^{4}),$
@@
and $F^{\alpha}=0$ with $\alpha=1,2$. Considering an invariant subset of this tensor under the combined $\operatorname{SO}(3)$ rotation (2.10) yields
@@
$B^{\alpha}$ $=(0,-\Psi^{4}),$
$C^{\alpha}$ $=(\Psi^{4},0),$
$D^{\alpha}$ $=(0,\Psi^{3}),$
$E^{\alpha}$ $=(-\Psi^{3},0),$
$F^{\alpha}$ $=(\Psi^{2},-\Psi^{1}),$ (2.28)
where this time we have chosen $A^{\alpha}=0$. It is again clear that only a subset of this tensor is an invariant under the transformation (2.10) and it is given by
@@ 2.1.2 Néel-type DMI
with $\nabla=(\partial_{1},\partial_{2},0)$ and $\bm{n}=(n^{1},n^{2},n^{3})$, which is due to the Rashba spin-orbit coupling (SOC), whereas the Bloch-type DMI (2.6) corresponds to the Dresselhaus SOC. Although, this DM term looks quite different from the Bloch type (2.6), the Néel-type DMI is simply obtained from the Bloch-type DMI by performing an $\operatorname{SO}(2)\subset\operatorname{SO}(3)$ rotation of the magnetization vector $(n^{1},n^{2},n^{3})\mapsto(n^{2},-n^{1},n^{3})$ and keeping the derivative vector fixed. The generalization of the above 2-dimensional DM term to a higher-dimensional DM term, however, is straightforward and somewhat different [^3] Often the opposite sign of the Néel-type DMI is used, which can be reached from the Bloch-type DMI simply by rotating the opposite way in the 1-2 plane: $(n^{1},n^{2},n^{3})\mapsto(-n^{2},n^{1},n^{3})$, or alternatively by flipping the sign of $\kappa$.:
@@
where now $\nabla=(\partial_{1},\partial_{2},\partial_{3},0)$ and $\bm{n}=(n^{1},n^{2},n^{3},n^{4})$. The symmetry properties are also quite clear and simple in this case, since the inner product is preserved by a combined rotation in $\rho\in\operatorname{SO}(3)\subset\operatorname{SO}(4)$ (see Eq. (2.14)) and $r\in\operatorname{SO}(3)$, with the symmetry-locking mechanism again given by $\rho=r$.
In fact, by a closer inspection it is clearly $\operatorname{SO}(3)$ invariant, as it is simply equal to the tensor (2.27) with $\Psi^{a}$ of Eq. (2.29), corresponding to the $\alpha$-part of the $\operatorname{SO}(3)$ invariant tensor (2.18). Indeed, a simple calculation shows that
$\frac{1}{2}\kappa\epsilon_{abcd}\epsilon^{abi4}\partial_{i}n^{c}n^{d}=\kappa(\delta_{c}^{i}\delta_{d}^{4}-\delta_{c}^{4}\delta_{d}^{4})\partial_{i}n^{c}n^{d}=\kappa(n^{4}\nabla\cdot\bm{n}-\bm{n}\cdot\nabla n^{4}),$ (2.32)

### v2

@@ Abstract
We propose a generalization of the theory of magnetic Skyrmions in chiral magnets in two dimensions to a higher-dimensional theory with magnetic Skyrmions in three dimensions and an $S^{3}$ target space, requiring a 4-dimensional magnetization vector. A physical realization of our theory could be made using a synthetic dimension, recently promoted and realized in condensed matter physics. In the simplest incarnation of the theory, we find a Skyrmion and a sphaleron – the latter being an unstable soliton. Including also the Skyrme term in the theory enriches the spectrum to a small metastable Skyrmion, an unstable sphaleron and a large stable Skyrmion.
@@ 1 Introduction
The magnetic Skyrmion is gaining recognition as a prominent soliton in 2-dimensional magnetic physics, particularly in the context of chiral magnets and thin films [1, 2, 3, 4, 5], as well as spintronics and quantum technologies [6, 7, 8, 9]. The name Skyrmion comes from the 3-dimensional soliton in the Skyrme model, proposed by T. Skyrme in the 1960s as a model of the nucleus [10, 11], further validated as a low-energy effective field theory of QCD at large $N_{c}$ (number of colors) by E. Witten [12, 13] and later used as a validation of the Sakai-Sugimoto model as evidence that the model really describes QCD at low energies [14]. The Skyrmion in 2 or 3 dimensions is a texture soliton, which differs from a defect-type soliton by mapping the entire 2- or 3-dimensional space to the target space. In particular, a magnetic Skyrmion in two dimensions is mapping the entire plane to an $S^{2}$ target space, with the latter being the normalized magnetization vector $\bm{n}=(n^{1},n^{2},n^{3})$ such that $\bm{n}\cdot\bm{n}=1$ [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. It differs drastically from the 3-dimensional Skyrmion by the fact that: 1) it is stabilized by a lower-dimensional operator (i.e. less derivatives than number of spatial dimensions) and 2) it is chiral. The latter implies that either (isolated) magnetic Skyrmions stably exist or the corresponding anti-Skyrmions, but not both of them. [^1] For isolated axially symmetric Skyrmions, this is the case and the stable (anti-)Skyrmions are those corresponding to a negative DMI energy (with its coefficient). However, both Skyrmions and anti-Skyrmion may exist for certain potentials and on domain lines, see e.g. [29, 23]. This is different from 3-dimensional Skyrmions, where a higher-dimensional operator (i.e. more derivatives than number of spatial dimensions) is used to stabilize the soliton – called a Skyrmion – against collapse. Indeed the higher-derivative (or smaller-derivative) terms are a way to avoid Derrick’s theorem [30].
@@
In this paper, we consider the generalization of the 2-dimensional magnetic Skyrmion with target space $S^{2}$ to a 3-dimensional magnetic Skyrmion [^2] A preliminary investigation started in Ref. [31]. with target space $S^{3}$.333 Solitons in 3He-B phase also have a 3-sphere for a target space as well as a spin-orbit interaction, but to the best of our knowledge does not give rise to a DMI term [32, 33]. Since the unit magnetization vector in three dimensions gives rise to $S^{2}$, we need a unit magnetization vector in four dimensions. [^4] We use the notation “magnetization vector” for the 4-dimensional generalization of the 3-dimension magnetization vector, although it is physically different. For example, it is not possible to write $\bm{B}\cdot\bm{n}$, since the 3-dimensional magnetic field has 3 components and a would-be 4-dimensional magnetic field cannot be dualized to a vector (the 4-dimensional Hodge dual of a 2-form is also a 2-form). We will henceforth denote the construct as a higher-dimensional magnetic Skyrmion.
@@
The physical realization of a 3-dimensional magnetic Skyrmion could be made by the incorporation of a synthetic (extra) dimension, as has been demonstrated in a number of physical systems – e.g. see Refs. [34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Until now, the 3-dimensional counterpart of 2D magnetic Skyrmions were magnetic Hopfions, 3-dimensional solitons living in a magnetization vector that also lives in three dimensions. Recent numerical computations [44, 45, 46, 47, 48] suggest that similar objects may exist at the nanoscale in frustrated and chiral magnets. Micrometer-sized Hopfions have recently been created and observed experimentally in chiral ferromagnetic fluids [49, 50]. The 3D magnetic Skyrmions which we propose, instead, have $S^{3}$ as a target space, the same as used in the standard Skyrme model in three dimensions.
@@
In this paper, we concentrate on the spherically symmetric case which enjoys $\operatorname{SO}(3)_{s}$ invariance, with the subscript $s$ denoting space. We take the generalized Dzyaloshinskii-Moriya interaction (DMI) [51, 52, 53] to have only one derivative of the magnetization vector and a single component of the same vector. In four dimensions, this means that such a term must be contracted with a tensor that possesses two $\operatorname{SO}(4)$-indices and a single spatial index: we denote such a tensor by $\Theta$. $\operatorname{SO}(4)$ invariance dictates that $\Theta$ is antisymmetric in $\operatorname{SO}(4)$-indices and we use this fact, as well as global rotations to rotate the tensor to its standard form. We further impose $\operatorname{SO}(3)_{\rm diag}$ invariance, where we are locking spatial rotations with an $\operatorname{SO}(3)_{t}\subset\operatorname{SO}(4)$ rotation, in order to obtain a spherically symmetric soliton (the subscript $t$ denotes the target space and ’diag’ denotes the locked combined symmetry). This procedure leads to two different invariant structures in the $\Theta$ tensor. However, only one of these two structures remains nonvanishing when a spherically symmetric Ansatz is imposed. We furthermore show that both the Bloch-type and the Néel-type DMI generalizations lead to the same term when $\operatorname{SO}(3)_{\rm diag}$ invariance is imposed.
@@ 2 Higher-dimensional magnetic Skyrmions
A possible extension of planar magnetic Skyrmions can be realized by increasing the dimension of the base space and target space from two to three dimensions [31]. This amounts to a magnetization unit vector that would live in a 4-dimensional Cartesian space. For a physical realization, advanced concepts such as synthetic dimensions must be considered, we will deal with this argument in Sec. 4. In symbols, the magnetization field is now a map $\bm{n}:\mathbb{R}^{3}\to S^{3}$ where $\bm{n}=(n^{1},n^{2},n^{3},n^{4})$ and $\bm{n}\cdot\bm{n}=1$. Let us consider the energy functional of the following form:
@@ 2.1.1 Bloch-type DMI
In order to put the higher-dimensional DM term on the most general standard form, we use $\operatorname{SO}(4)$ and $\operatorname{SO}(3)_{s}$ transformations to simplify the tensor $\Theta$ as much as possible. Since $\Theta$ is contracted with the $\operatorname{SO}(4)$-invariant tensor $\epsilon_{abcd}$, we can from here on assume that it is antisymmetric in the $\operatorname{SO}(4)$-indices $a$ and $b$:
@@
Performing simultaneously both orthogonal transformations, $R\in\operatorname{SO}(4)$ and $r\in\operatorname{SO}(3)_{s}$, we get
@@
which is invariant only under the $\operatorname{SO}(3)_{t}$ subgroup represented by $\rho\in\operatorname{SO}(3)_{t}\subset\operatorname{SO}(4)$ such that
@@
the rotation $\rho$ of Eq. (2.14) can only be used to diagonalize the submatrix consisting of $A^{i}$, $B^{i}$ and $C^{i}$. Since the submatrix is $3\times 3$, it only possesses 1 purely imaginary pair of nonvanishing eigenvalues, as well as a zero eigenvalue. There is still no reason for the $\operatorname{SO}(3)_{s}$ vectors to be aligned, so we can only diagonalize one of the $\Theta^{i}$ submatrices, say the $i=3$ again. We thus arrive at
$\Theta^{\mu}=\frac{1}{2}\begin{pmatrix}0&A^{\mu}&B^{\mu}&D^{\mu}\\ -A^{\mu}&0&C^{\mu}&E^{\mu}\\ -B^{\mu}&-C^{\mu}&0&F^{\mu}\\ -D^{\mu}&-E^{\mu}&-F^{\mu}&0\end{pmatrix},\qquad\Theta^{3}=\frac{1}{2}\begin{pmatrix}0&A^{3}&0&D^{3}\\ -A^{3}&0&0&E^{3}\\ 0&0&0&F^{3}\\ -D^{3}&-E^{3}&-F^{3}&0\end{pmatrix},$ (2.16)
with $\mu=1,2$. Using the $r\in\operatorname{SO}(3)_{s}$ degrees of freedom, we can eliminate one symbol in $\Theta^{\mu}$, by rotating it into the $i=3$ direction; we may choose this to be $F^{\mu}=0$. We will denote this the standard form of the $\Theta$ tensor. One result of this computation is that the minimal degrees of freedom of the $\Theta$ tensor are 14.
At this point, it will be useful to identify a subset of the $\Theta$ tensor in standard form, that is invariant under a combined $\operatorname{SO}(3)_{t}\subset\operatorname{SO}(4)$ and $\operatorname{SO}(3)_{s}$ symmetry. The transformation is again that of Eq. (2.10) with $R$ of Eq. (2.14), it takes the form
$\Theta^{i}\to\frac{1}{2}\begin{pmatrix}\rho^{\rm T}\begin{pmatrix}0&A^{j}&B^{j}\\ -A^{j}&0&C^{j}\\ -B^{j}&-C^{j}&0\end{pmatrix}\rho&\rho^{\rm T}\begin{pmatrix}D^{j}\\ E^{j}\\ F^{j}\end{pmatrix}\\ \begin{pmatrix}-D^{j}&-E^{j}&-F^{j}\end{pmatrix}\rho&0\end{pmatrix}r_{j}^{\phantom{j}i},$ (2.17)
where $\rho\in\operatorname{SO}(3)_{t}$ and $r\in\operatorname{SO}(3)_{s}$. From the above equation, it is clear that invariance requires the triplet $A,B,C$ to transform under the adjoint representation of $\operatorname{SO}(3)$ and the triplet $D,E,F$ to transform under the fundamental representation. This means that invariance under the combined rotation with the locking of the two $\operatorname{SO}(3)$’s according to Eq. (2.14) with $\rho=r$, is attained by the tensor
@@
$B^{\mu}$ $=(-\Gamma^{3},0),$
$C^{\mu}$ $=(0,-\Gamma^{3}),$
$D^{\mu}$ $=(-\Gamma^{4},0),$
$E^{\mu}$ $=(0,-\Gamma^{4}),$
@@
and $F^{\mu}=0$ with $\mu=1,2$. Considering an invariant subset of this tensor under the combined $\operatorname{SO}(4)$ and $\operatorname{SO}(3)_{s}$ rotation (2.10) yields
@@
$B^{\mu}$ $=(0,-\Psi^{4}),$
$C^{\mu}$ $=(\Psi^{4},0),$
$D^{\mu}$ $=(0,\Psi^{3}),$
$E^{\mu}$ $=(-\Psi^{3},0),$
$F^{\mu}$ $=(\Psi^{2},-\Psi^{1}),$ (2.28)
where this time we have chosen $A^{\mu}=0$. It is again clear that only a subset of this tensor is an invariant under the transformation (2.10) and it is given by
@@ 2.1.2 Néel-type DMI
with $\nabla=(\partial_{1},\partial_{2},0)$ and $\bm{n}=(n^{1},n^{2},n^{3})$, which is due to the Rashba spin-orbit coupling (SOC), whereas the Bloch-type DMI (2.6) corresponds to the Dresselhaus SOC. Although, this DM term looks quite different from the Bloch type (2.6), the Néel-type DMI is simply obtained from the Bloch-type DMI by performing an $\operatorname{SO}(2)_{t}\subset\operatorname{SO}(3)_{t}$ rotation of the magnetization vector $(n^{1},n^{2},n^{3})\mapsto(n^{2},-n^{1},n^{3})$ and keeping the derivative vector fixed. The generalization of the above 2-dimensional DM term to a higher-dimensional DM term, however, is straightforward and somewhat different [^5] Often the opposite sign of the Néel-type DMI is used, which can be reached from the Bloch-type DMI simply by rotating the opposite way in the 1-2 plane: $(n^{1},n^{2},n^{3})\mapsto(-n^{2},n^{1},n^{3})$, or alternatively by flipping the sign of $\kappa$.:
@@
where now $\nabla=(\partial_{1},\partial_{2},\partial_{3},0)$ and $\bm{n}=(n^{1},n^{2},n^{3},n^{4})$. The symmetry properties are also quite clear and simple in this case, since the inner product is preserved by a combined rotation in $\rho\in\operatorname{SO}(3)_{t}\subset\operatorname{SO}(4)$ (see Eq. (2.14)) and $r\in\operatorname{SO}(3)_{s}$, with the symmetry-locking mechanism again given by $\rho=r$.
In fact, by a closer inspection it is clearly $\operatorname{SO}(3)_{\rm diag}$ invariant, as it is simply equal to the tensor (2.27) with $\Psi^{a}$ of Eq. (2.29), corresponding to the $\alpha$-part of the $\operatorname{SO}(3)_{\rm diag}$ invariant tensor (2.18). Indeed, a simple calculation shows that
$\frac{1}{2}\kappa\epsilon_{abcd}\epsilon^{abi4}\partial_{i}n^{c}n^{d}=\kappa(\delta_{c}^{i}\delta_{d}^{4}-\delta_{c}^{4}\delta_{d}^{i})\partial_{i}n^{c}n^{d}=\kappa(n^{4}\nabla\cdot\bm{n}-\bm{n}\cdot\nabla n^{4}),$ (2.32)